I need an linear equation that is perpendicular to the line y=4x+5 and that passes through the point (0,-3)??: I need an linear equation that is perpendicular to the line y=4x+5 and that passes through the point (0,-3)??
Answer by nerdybill(1280)

I need an linear equation that is perpendicular to the line y=4x+5 and that passes through the point (0,-3)??
.
From inspecting:
y=4x+5
we see that the slope is 4
.
A line that is perpendicular is a negative reciprocal:
-1/4
.
Now, we have a slope (-1/4) and a point (0, -3)
Plug it into the "point slope form":
y - y1 = m(x-x1)
y - (-3) = (-1/4)(x-0)
y + 3 = (-1/4)x
4y + 12 = -x
4y = -x - 12
y = (-1/4)x - 3
y = -x/4 - 3 (This is what they're looking for)